A boat traveled 16 miles down a river and then returned. The total time for the round trip was 6 hours, and the rate of the boat instill water is 6 mph. Find the rate of the current.
Answer by mananth(16949)

A boat traveled 16 miles down a river and then returned. The total time for the round trip was 6 hours, and the rate of the boat instill water is 6 mph. Find the rate of the current.
Let the speed of the current be x
down the river its speed will be 6+x . Time taken = 16/ 6-x hours
against the flow the speed will be 6-x. Time taken = 16/ 6-x hours
Time taken down the river + time taken up the river = 6
16/ 6-x + 16/ 6+x = 6
96-16x+96+16x / (6+x)(6-x) = 6
192 = 6* (36-x^2)
192=216 - 6x^2
192-216= 6x^2
-24 = 6x^2
x^2 -= -24/-6
x^2=4
x=2 the rate of the current